A 5.00-kg block is pulled a distance of 1.50 m along a frictionless horizontal surface by an applied force of 30.0 N at an angle of 30.0 above the horizontal. (a) What is the normal force exerted by the surface on the block, assuming the block never loses contact with the surface? (b) What is the acceleration of the block? (c) What is the final speed of the block?


1
Expert's answer
2021-03-10T17:16:53-0500

a)


"N=mg-F\sin{30}=(5)(9.8)-30\sin{30}=34\ N"

b)


"5a=30\sin{30}\\a=3\frac{m}{s^2}"

c)


"v^2=2as\\v^2=2(3)(1.5)\\v=3\frac{m}{s}"